# The fate of entanglement

### v1

@@ Abstract
[…]
${}^{1}$épartement de Physique, Université de Montréal, Montréal, QC H3C 3J7, Canada
${}^{2}$entre de Recherches Mathématiques,
@@
${}^{3}$nstitut Courtois, Université de Montréal, Montréal, QC H2V 0B3, Canada
I A. Entanglement measures and witness
We give the definitions of the entanglement-related quantities we discuss in the main text, namely the logarithmic negativity $\mathcal{E}$, the geometric entanglement $\mathcal{D}$ and the witness $W$.
@@ I.1 Logarithmic negativity
We consider the density matrix $\rho$ matrix pertaining to two subsystems 1 and 2. The logarithmic negativity VW02 ; plenio2005logarithmic is
@@
where $\lVert X\rVert=\operatorname{Tr}\sqrt{XX^{\dagger}}$ is the trace-norm, and $\rho^{T_{1}}$ is the partially-transposed density matrix with respect to subsystem 1. This entanglement measure is related to the Peres separability criterion for density matrices peres1996separability . In full generality, a vanishing logarithmic negativity is only a necessary condition for separability, but in the case of two spins (or qubits), it is both necessary and sufficient HORODECKI19961 .
@@ I.2 Geometric entanglement
The geometric entanglement vedral1997quantifying is the distance between the state $\rho$ and the closest separable state,
@@
I.3 Genuine 3-party witness
The witness $W$ is defined from the matrix elements $\rho_{ij}$, $i,j=1,\dots,8,$ of a 3-spin density matrix. It reads guhne2010separability
@@ I.3 Genuine 3-party witness
Positive $W>0$ indicates the presence of genuine 3-party entanglement, whereas for $W\leqslant 0$ a definitive conclusion cannot be made. We thus define $W=0$ in this case. The witness $W$ is basis-dependent. Numerically, we thus maximize its value over all possible local unitary transformations $(U_{1}\otimes U_{2}\otimes U_{3})\rho(U_{1}^{\dagger}\otimes U_{2}^{\dagger}\otimes U_{3}^{\dagger})$, where $U_{j}$ is a generic $2\times 2$ unitary matrix for the spin $j$. While it is not an entanglement quantifier, large values of $W$ generically correspond to stronger genuine 3-party entanglement. For instance, the witness is maximal for Werner states, which is a maximally-entangled 3-qubit state, and is minimal for the maximally mixed state (the identity).
II B. Phase diagram of the molecular quantum magnet

### v2

@@ Abstract
Gilles Parez1,2 and William Witczak-Krempa1,2,3
1Département de Physique, Université de Montréal, Montréal, QC H3C 3J7, Canada
2Centre de Recherches Mathématiques,
@@
3Institut Courtois, Université de Montréal, Montréal, QC H2V 0B3, Canada
I A. Entanglement measures and criteria
We give the definitions of the entanglement-related quantities we discuss in the main text, namely the logarithmic negativity $\mathcal{E}$, the geometric entanglement $\mathcal{D}$ and the genuine 3-party entanglement criterion $W$.
@@ I.1 Logarithmic negativity
We consider the density matrix $\rho$ matrix pertaining to two subsystems 1 and 2. The logarithmic negativity [30, 31] is
@@
where $\lVert X\rVert=\operatorname{Tr}\sqrt{XX^{\dagger}}$ is the trace-norm, and $\rho^{T_{1}}$ is the partially-transposed density matrix with respect to subsystem 1. This entanglement measure is related to the Peres separability criterion for density matrices [43]. In full generality, a vanishing logarithmic negativity is only a necessary condition for separability, but in the case of two spins (or qubits), it is both necessary and sufficient [32].
@@ I.2 Geometric entanglement
The geometric entanglement [3] is the distance between the state $\rho$ and the closest separable state,
@@
I.3 Genuine 3-party entanglement criterion
The criterion $W$ is defined from the matrix elements $\rho_{ij}$, $i,j=1,\dots,8,$ of a 3-spin density matrix. It reads [33]
@@ I.3 Genuine 3-party entanglement criterion
Positive $W>0$ indicates the presence of genuine 3-party entanglement, whereas for $W\leqslant 0$ a definitive conclusion cannot be made. We thus define $W=0$ in this case. The criterion $W$ is basis-dependent. Numerically, we thus maximize its value over all possible local unitary transformations $(U_{1}\otimes U_{2}\otimes U_{3})\rho(U_{1}^{\dagger}\otimes U_{2}^{\dagger}\otimes U_{3}^{\dagger})$, where $U_{j}$ is a generic $2\times 2$ unitary matrix for the spin $j$. While it is not an entanglement quantifier, large values of $W$ generically correspond to stronger genuine 3-party entanglement. For instance, $W$ is maximal for Werner states, which are maximally-entangled 3-qubit states, and is minimal for the maximally mixed state (the identity).
II B. Test of the full-rank hypothesis
We test the full-rank hypothesis (FRH) for the icosahedral molecule at $h=3$ for three adjacent sites. In Fig. S1 we plot the minimal eigenvalues $\lambda_{\rm min}$ of the 3-spin reduced density matrices for each eigenstate of the Hamiltonian, labelled by the corresponding energy $E$. All the minimal eigenvalues are strictly positive, which indicates that the FRH is satisfied. The smallest $\lambda_{\rm min}$ occurs in the groundstate.
Figure S1: Test of the full-rank hypothesis. Minimal eigenvalue $\lambda_{\rm min}$ of the reduced density matrices for three adjacent spins in the icosahedral molecule at $h=3$ for the whole spectrum, labelled by the corresponding energy $E$.
III C. Phase diagram of the molecular quantum magnet
@@
IV D. Details on the quench protocol
In the quench protocol, we initialise the system in a pure product state of up and down spins, and let it evolve under the Ising Hamiltonian discussed above with $h=3$. We illustrate the initial state in Fig. S3. This is a planar representation of the icosahedral molecule, where blue sites are initialised in a spin-up state, whereas orange one are initialised in a spin-down state. The coloured face represents the 3-spin subsystem for which we compute $\mathcal{D}$ and $W$ during the time evolution.
Figure S3: Initial state in the quench protocol. Blue sites are initialised in a spin-up state, whereas orange one are initialised in a spin-down state. We focus on the coloured face during the time evolution.